OpenStudy (tinylittlehelper):

My mean absolute deviation (MAD) calculator isn't working correctly; it should be saying 1.6, not 6.4...

OpenStudy (tinylittlehelper):

//place values here var a = 10, b = 11, c = 11, d = 12, e = 16, f = 0, g = 0; //place number of values here var numOfValues = 5; //DO NOT CHANGE OR U WILL GET IT WRONG! var sum = a + b + c + d + e +f + g; var mean = sum / numOfValues; var aa = Math.abs(mean - a), bb = Math.abs(mean - b), cc = Math.abs(mean - c),dd = Math.abs(mean - d),ee = Math.abs(mean - e),ff = Math.abs(mean - f),gg = Math.abs(mean - g); var sumTwo = aa + bb + cc + dd + ee + ff + gg; var meanTwo = sumTwo / numOfValues; var answer = meanTwo; background(112, 79, 79); text(answer,20,20);

OpenStudy (lyrae):

``` var numOfValues = 5 ``` should be ``` var numOfValues = 7 ``` ?